# references.bib A comparison of methods for estimating the average treatment effect on the treated for externally controlled trials

### v1

@@ 4.1.2 Nonmatching-Based Methods
The bias of AIPW_SL is much smaller than the AIPW when the nonlinear effect is present and comparable to the AIPW in other settings. Attributed to the doubly robustness and SuperLearner’s flexibility, the AIPW_SL always presents a small bias in all of our settings. In addition, the SD of AIPW_SL is also very small for all settings in Scenarios 1 and 2. Therefore, the MSE of AIPW_SL is always one of the smallest, if not the smallest, of all settings for Scenarios 1 and 2. Another doubly robust estimator with the SuperLearner, TMLE_SL, has similar bias/SD/MSE as the AIPW_SL in Scenario 1, but higher SD and MSE in Scenario 2.
@@
Under Scenario 3 where an important confounder is not considered in estimating ATT, a substantial bias is introduced in all models. Therefore, the bias due to the violation of unconfoundedness cannot be mitigated by model selection. Although the unconfoundedness assumption cannot be formally examined, sensitivity analyses have been developed to assess whether a significant result is sensitive to potential unmeasured confounding. For example, a popular sensitivity analysis is based on the E-value measuring the minimum strength of association that an unmeasured confounder would need to have with both the treatment and outcome to explain away the observed causal effect [vanderweele2017sensitivity]. It is important to note, however, that the application of the E-value is constrained by certain limitations, and its interpretation should be approached cautiously [blum2020use].
@@ 4.2 Type I error control
In terms of type I error control, the PSM and PSM_1:2 models perform best for all settings in Scenarios 1 and 2 because of their small biases and accurate estimation of SDs. The MDM model also largely controls type I errors across all settings in Scenarios 1 and 2.
@@
This study compared the performance of various methods for estimating the ATT in single-arm clinical trials with external RWD controls. Through simulations under different identifiability assumptions, we evaluated evaluation metrics of 10 causal inference methods, including matching-based and nonmatching-based approaches. We found that the doubly robust methods generally outperformed other approaches in terms of bias. However, no single method consistently outperformed others in terms of MSE. Additionally, the identifiability assumptions played a critical role in the models’ performance, with violations of the positivity and unconfoundedness assumptions leading to potential inflation of type I error and bias, respectively. Among the matching-based methods, PSM and PSM_1:2 showed robustness under various settings, but their performance was affected by violations of the positivity assumption and interactions between treatment variables and covariates. By contrast, nonmatching methods in general had smaller SDs than matching-based methods due to their ability to utilize more data. AIPW_SL and TMLE_SL exhibited advantages in terms of bias and MSE due to their double robustness and flexibility, but they tended to underestimate the SD and experienced type I error inflation with violated positivity assumption. In summary, the selection of the appropriate method for causal inference in single-arm trials with external RWD controls should carefully consider identifiability assumptions and potential biases in realistic situations.
According to findings in this study, we would like to recommend the following procedures for planning the single-arm trial with external control.
@@
Consider the unconfoundedness assumption in the design phase: The unconfoundedness assumption can be met if all confounders are identified and controlled. However, if an important confounding factor is not taken into account, the analysis results can be highly biased. Every effort should be made to identify all important confounders through the literature and expert knowledge. In addition, it should be ensured that the information from all important confounders is measured in the planned single-arm test and is available in the RWD used for the ECA. Furthermore, sensitivity analyses such as the E-value or other tipping point analyses should be pre-planned during the design stage.
@@
Select appropriate analytic models for the primary and sensitivity analyses: One of the most important requirements for the primary analysis method is to ensure the control of type I errors. After the type I error control is taken into account, the accuracy of the causal effect estimation is the next important consideration. For trials in which the positive assumption cannot be validly justified in the design phase, we suggest using methods that can largely control type I errors even if the positive assumption does not hold, e.g., the PSM or MDM model. When the expected prevalence rate of treatment is low, then 1:n with n2 matching should be considered. For experiments where the positivity assumption can be validly justified, a doubly robust approach such as TMLE_SL or AIPW_SL may be recommended as the primary analysis method to obtain better estimation precision.
@@
Evaluate the positivity assumption: The plot of the PS distribution should be plotted to evaluate the positivity assumption. If the positivity assumption is violated, then results based on methods that are strongly influenced by the violation of this assumption, such as doubly robust methods with the SuperLearner, are not reliable.
@@
Conduct sensitivity analyses for the unconfoundedness assumption: Although the unconfoundedness assumption in general cannot be tested directly, sensitivity analyses should be performed to assess whether the results are robust against the unmeasured confounder.
Acknowledgement

### v2

@@ 4.1.2 Nonmatching-Based Methods
The bias of AIPW_SL is much smaller than the AIPW when the nonlinear effect is present and comparable to the AIPW in other settings. Attributed to the doubly robustness and SuperLearner’s flexibility, the AIPW_SL always presents a small bias in all of our settings in Scenarios 1 and 2. In addition, the SD of AIPW_SL is also very small for all settings in Scenarios 1 and 2. Therefore, the MSE of AIPW_SL is always one of the smallest, if not the smallest, of all settings for Scenarios 1 and 2. Another doubly robust estimator with the SuperLearner, TMLE_SL, has similar bias/SD/MSE as the AIPW_SL in Scenario 1, but higher SD and MSE in Scenario 2.
@@
Under Scenario 3 where an important confounder is not considered in estimating ATT, a substantial bias is introduced in all models. Therefore, the bias due to the violation of unconfoundedness cannot be mitigated by model selection. Although the unconfoundedness assumption cannot be formally examined, sensitivity analyses have been developed to assess whether a significant result is sensitive to potential unmeasured confounding. For example, a popular sensitivity analysis is based on the E-value measuring the minimum strength of association that an unmeasured confounder would need to have with both the treatment and outcome to explain away the observed causal effect [vanderweele2017sensitivity]. It is important to note, however, that the application of the E-value is constrained by certain limitations, and its interpretation should be approached with caution [blum2020use].
@@ 4.2 Type I error control
In terms of type I error control, the PSM, PSM_1:2, and MDM models perform best for all settings in Scenarios 1 and 2 because of their small biases and accurate estimation of SDs.
@@
4.3 Scope and limitations of simulation findings
We would like to emphasize that the primary purpose of our simulations is to identify scenarios and settings that reveal the limitations of the causal inference methods, rather than to provide a comprehensive evaluation under all possible conditions, which is not feasible. As a result, the comparison results for performance of estimation and type I error control may not directly generalize to other simulation settings. When considering specific methods, we recommend that trialists perform thorough evaluations tailored to their particular needs prior to implementation.
@@ 4.3 Scope and limitations of simulation findings
This study compared the performance of various methods for estimating the ATT in single-arm clinical trials with external RWD controls. Through simulations under different identifiability assumptions, we evaluated metrics of 10 causal inference methods, including matching-based and nonmatching-based approaches. We found that the doubly robust methods generally outperformed other approaches in terms of bias. However, no single method consistently outperformed others in terms of MSE. Additionally, the identifiability assumptions played a critical role in the models’ performance, with violations of the positivity and unconfoundedness assumptions leading to potential inflation of type I errors and biases. Among the matching-based methods, PSM, PSM_1:2, and MDM showed robustness under various settings, but their performance was affected by the violation of the positivity assumption and the interaction between the treatment variable and the covariate. By contrast, nonmatching methods in general had smaller SDs than matching-based methods due to their ability to utilize more data. AIPW_SL and TMLE_SL exhibited advantages in terms of bias and MSE due to their double robustness and flexibility, but they tended to underestimate the SD and experienced type I error inflation with violated positivity assumption.
In light of these findings, it is essential to recognize the limitations of causal inference methods for single-arm trials with external controls, especially when any identifiability assumption can be violated. Violations of these assumptions can lead to biased estimates of treatment effects and an increased risk of type I errors. Therefore, it is crucial to have careful considerations of these factors in both the design and analysis phases of such trials. Accordingly, we would like to recommend the following procedures for planning the single-arm trial with an external control.
@@
Consider the unconfoundedness assumption in the design phase: The unconfoundedness assumption can be met if all confounders are identified and controlled. However, if an important confounding factor is not taken into account, the analysis results can be highly biased. Every effort should be made to identify all important confounders through the literature and expert knowledge. In addition, it should be ensured that the information from all important confounders is measured in the planned single-arm trial and is available in the RWD used for the ECA. Furthermore, sensitivity analyses such as the E-value or other tipping point analyses should be pre-planned during the design stage.
@@
Select appropriate analytic models for the primary and sensitivity analyses: One of the most important requirements for the primary analysis method is to ensure the control of the type I error. After the type I error control is taken into account, the accuracy of the causal effect estimation is the next important consideration. For trials in which the positive assumption cannot be validly justified in the design phase, we suggest using methods that can largely control the type I error even if the positive assumption does not hold, e.g., the PSM or MDM model. When the expected prevalence rate of treatment is low, then 1 to r matching with r 2 should be considered. For experiments where the positivity assumption can be validly justified, a doubly robust approach such as TMLE_SL or AIPW_SL may be recommended as the primary analysis method to obtain better estimation precision.
@@
Evaluate the positivity assumption: The plot of the PS distribution should be plotted to evaluate the positivity assumption. If the positivity assumption appears to be clearly violated, then results based on methods that are strongly influenced by the violation of this assumption, such as doubly robust methods with the SuperLearner, may not be reliable and should be interpreted with caution.
@@
Conduct sensitivity analyses for the unconfoundedness assumption: Although the unconfoundedness assumption in general cannot be tested directly, sensitivity analyses should be performed to assess whether the primary analysis results are robust against the unmeasured confounder.
Acknowledgements
@@
Supplementary materials
Supplementary materials are available online with this paper.